OpenStudy (anonymous):

The function C(n) = 1.25n represents the cost, C(n), of n sodas. What is the cost of purchasing 12 sodas?

OpenStudy (ddcamp):

Just replace n in the equation with the number of sodas.

OpenStudy (anonymous):

which would be ? its easier said then done ?

OpenStudy (ddcamp):

The equation is C(n) = 1.25×n What would C(n) be if n was 12? C(12) = 1.25×12

OpenStudy (anonymous):

1.25*12=15